# Next-Generation Database Interfaces: A Survey of LLM-based Text-to-SQL

### v3

@@ IV-A In-context Learning
The impact of cross-domain few-shot examples is also being studied [54]. When incorporating in-domain and out-of-domain examples with different numbers, the in-domain demonstration outperforms zero-shot and out-of-domain examples, which will also perform better as the number of examples rises. To explore the detailed construction of input prompt, [53] compare the concise and verbose prompt design approaches. The former style splits the schema, the column names, and the primary and foreign keys by bar, and the latter organizes them as natural language descriptions.
$\mathbf{C}_{1}$-Decomposition
As an intuitive solution, decomposing a challenging user question into simpler sub-questions or using multi-components for implementation can reduce the complexity of the overall text-to-SQL task [51, 8]. Dealing with less complexity, LLMs have the potential to elicit more accurate SQL generation. The decomposition-based approaches for LLM-based text-to-SQL are categorized into two paradigms: (1) sub-task decomposition, provides additional parsing to assist the final SQL generation by decomposing the overall text-to-SQL task into more manageable effective sub-tasks (e.g., schema linking [71], domain classification [54]). (2) sub-question decomposition, divides the user question into sub-questions to reduce the question’s complexity and difficulty, then generates the sub-SQL by solving these questions to deduce the final SQL query. The technical novelty of the decomposition paradigm
DIN-SQL [8] proposed a decomposed in-context learning method consisting of four modules: schema linking, classification & decomposition, SQL generation, and self-correction. DIN-SQL first generates the schema linking between the user question and the target database; the following module decomposes the user question into correlated sub-questions and does a difficulty classification. Based on the above information, the SQL generation module generates a corresponding SQL, and the self-correction module identifies and corrects the potential errors in the predicted SQL. This approach comprehensively considers the sub-questions decomposition as a module of sub-tasks decomposition. Coder-Reviewer [56] framework proposed a re-ranking method, combining Coder models for the generation and Reviewer models to evaluate the likelihood of the instruction. Refer to the Chain-of-Thought [103] and Least-to-Most prompting [104], QDecomp [51] introduce question decomposition prompting, which follows the question reduction stage in least-to-most prompting and instruct the LLM to decompose the original complex question as the intermediate reasoning steps. C3 [30] consists of three key components: clear prompting, calibration bias prompting, and consistency; these components are accomplished by assigning ChatGPT with different tasks. Firstly, the clear prompting component generates the schema linking and the distilled question-relevant schema as a clear prompt. Then, a multi-turn dialogue about text-to-SQL hints is utilized as a calibration bias prompt, which combines with the clear prompt to guide the SQL generation. The generated SQL queries are selected by consistency and execution-based voting to get the final SQL. MAC-SQL [57] presents a multi-agent collaborating framework; the text-to-SQL process is finished as the collaboration of the agents: Selector, Decomposer, and Refiner. The Selector preserves relevant tables for user questions; the Decomposer breaks down user questions into sub-questions and provides solutions; finally, the Refiner validates and refines the defective SQL. DEA-SQL [58] introduces a workflow paradigm aiming to enhance the attention and problem-solving scope of LLM-based text-to-SQL through decomposition. This method decomposes the overall task, enabling the SQL generation module to have the corresponding prerequisite (information determination, question classification) and subsequent (self-correction, active learning) sub-tasks. The workflow paradigm allows the LLM to generate more accurate SQL queries. SGU-SQL [32] is a structure-to-SQL framework, leveraging the inherent structure information to assist SQL generation. Specifically, the framework constructs a graph structure for the user question and the corresponding database respectively, then uses the encoded graphs to construct structure linking [105, 106]. A meta operator decomposes the user question with a grammar tree and finally designs the input prompt with meta-operation in SQL. MetaSQL [59] introduces a three-stage approach for SQL generation: decomposition, generation, and ranking. The decomposition stage uses semantic decomposition and metadata composition to process the user question. Taking the previously processed data as input, a text-to-SQL model using metadata-conditioned generation to generate some candidate SQL queries. Finally, a two-stage ranking pipeline is applied to get a global-optimal SQL query. PET-SQL [60] proposed a prompt-enhanced two-stage framework. Firstly, an elaborated prompt instructs the LLMs to generate preliminary SQL (PreSQL) where some few-shot demonstrations are selected based on similarity. Then, schema linking is found based on PreSQL and combined to prompt the LLMs to generate the Final SQL (FinSQL). Finally, multiple LLMs are utilized to generate a FinSQL, ensuring consistency based on the execution results.
$\mathbf{C}_{2}$-Prompt Optimization
As previously introduced, few-shot learning is widely studied for prompting LLMs [85]. For LLM-based text-to-SQL with in-context learning, trivial few-shot approaches obtained promising results [33, 8, 9], further optimization of few-shot prompting has the potential to elicit better performance. Since the accuracy of SQL generation in off-the-shelf LLMs largely depends on the quality of the corresponding input prompt [107], many decisive factors that can influence the quality of the prompt have become points of focus of the current research [9] (e.g., quality and quantity in the few-shot organization, the similarity between user questions and few-shots instances, external knowledge/hints) The process of improving prompt quality can be summarized as the prompt’s optimization, including advanced few-shot sampling strategies, schema information augmentation, and external knowledge integration.
DESEM [62] is a prompt engineering framework with de-semanticization and skeleton retrieval. The framework first employs domain-specific words masking module to remove the semantic tokens in user questions that preserve the intentions. It then utilizes an adjustable prompting module that retrieves the few-shot examples with identical question intentions and incorporates schema-relevance filtering to guide the LLM’s SQL generation. The QDecomp [51] framework introduces the InterCOL mechanism to incrementally incorporate the decomposed sub-questions with correlative table and column names. With difficulty-based selection, the few-shot examples for QDecomp are difficult-level sampled. Besides similarity-diversity sampling, [52] proposed SD+SA+Voting (Similarity-Diversity+Schema Augmentation+Voting) sampling strategy. They first employ semantic similarity and $k$-Means cluster diversity for sampling few-shot examples and then enhance the prompt with schema knowledge (semantic or structure augmentation). C3 [30] framework comprises a clear prompting component, which takes the question and schema as the LLMs input, generates a clear prompt that includes a schema that removes the redundant information irrelevant to the user question and a schema linking, and also a calibration component providing hints. The LLMs take their composition as context-augmented prompts for SQL generation. A retrieval-augmented framework is introduced with sample-aware prompting [64], which simplifies the original question and extracts the question skeleton from the simplified question, then finishes the sample retrieval in the repository according to skeleton similarities. The retrieved samples are combined with the original question for few-shot prompting. ODIS [54] introduces the selection of a sample with out-of-domain demonstrations and in-domain synthetic data, which retrieves few-shot demonstrations from hybrid sources to augment the prompt representations. DAIL-SQL [9] proposed a novel approach to address the issues in few-shot sampling and organization, presenting a better balance between the quality and quantity of few-shot examples. DAIL Selection first masks domain-specific words in user and few-shot example questions, then ranks the candidate examples based on the embedded Euclidean distance. Meanwhile, the similarity between the pre-predicted SQL queries is calculated. Finally, the selection mechanism obtains the similarity-sorted candidates according to the pre-set criteria. The few-shot examples are guaranteed good similarity with both questions and SQL queries with this method. ACT-SQL [49] proposed dynamic examples in few-shot prompting, which is selected according to similarity score. FUSED [65] are presented to build a high-diversity demonstrations pool through human-free multiple-iteration synthesis to improve the diversity of the few-shot demonstrations. The pipeline of FUSED samples the demonstrations to be fused by clustering, then fuse the sampled demonstrations to construct the pool to enhance few-shot learning. Knowledge-to-SQL [31] framework aims to build a Data Expert LLM (DELLM) to provide knowledge for SQL generation. The DELLM is trained by supervised fine-tuning using human expert annotations [33] and further refined by preference learning with the database’s feedback. DELLM generates four categories of knowledge, the well-designed methods (e.g. DAIL-SQL [9], MAC-SQL [57]) incorporating the generated knowledge to achieve better performance for LLM-based text-to-SQL with in-context learning.
$\mathbf{C}_{3}$-Reasoning Enhancement
LLMs have exhibited promising capabilities in tasks involving commonsense reasoning, symbolic reasoning, and arithmetic reasoning [108]. For text-to-SQL tasks, numeric and synonym reasoning frequently occur in realistic scenarios [33, 41]. The prompting strategies for LLMs’ reasoning have the potential to enhance SQL generation capabilities. Recent studies primarily focus on integrating well-designed reasoning-enhanced methods for text-to-SQL adaptation, improving LLMs to address the challenge about complex questions that require sophisticated reasoning [^3] The difference between multi-step reasoning (e.g., Chain of Thought) and the decomposition paradigm is that the former research focuses on advancing inherent reasoning within a single-turn generation, whereas the latter study involves using different components to assist the final generation through multiple calls of LLMs. and the self-consistency in SQL generation.
Chain-of-Thoughts (CoT) prompting technique [103] involves a comprehensive reasoning process that guides LLMs towards accurate deduction, eliciting reasoning in LLMs. The study of LLM-based text-to-SQL utilizes CoT prompting as rule implication [9], which setting the instruction “Let’s think step by step” in prompt construction [33, 51, 9, 32]. However, the straightforward (original) CoT strategy has not demonstrated the potential in text-to-SQL tasks that it has in other reasoning tasks; studying CoT for adaptations is still an ongoing research [51]. Since CoT prompting always uses static examples with human annotation for demonstrations, which requires empirical judgment for the effective selection of few-shot examples, and manual annotating is also an essential need. As a solution, ACT-SQL [49] proposed a method to generate CoT examples automatically. Specifically, given a question, ACT-SQL truncates a set of slices of the question and then enumerates every column appearing in the corresponding SQL query. Each column will be linked with its most relevant slice through the similarity function and appended to the CoT prompt. Through systematical study for enhancing LLMs SQL generation incorporating CoT prompting, QDecomp [51] presents a novel framework to address the challenge for CoT how to come up with the reasoning steps to predict the SQL query. The framework utilizes every slice of the SQL query to construct a logical step in CoT reasoning, then employs natural language templates to articulate each slice of the SQL query and arranges them in the logical execution order. Least-to-Most [104] is another prompting technique that decomposes questions into sub-questions and then sequentially solves them. As iterative prompting, pilot experiments [51] demonstrate that it may be unnecessary for text-to-SQL parsing. Using detailed reasoning steps tends to have more error propagation issues. As a variant of CoT, Program-of-Thoughts (PoT) prompting strategy [109] are proposed to enhance arithmetic reasoning for LLMs. Through evaluation [55], PoT enhances the LLM for SQL generation, especially in complicated datasets [33]. SQL-CRAFT [55] are proposed to enhance LLM-based SQL generation, which incorporates PoT prompting for Python-enhanced reasoning. PoT strategy requires the model to simultaneously generate the Python code and the SQL queries, enforcing the model to incorporate Python code in its reasoning process. Self-Consistency [110] is a prompting strategy improving reasoning in LLMs, which leverages the intuition that a complex reasoning problem typically admits multiple different ways of thinking, leading to its unique correct answer. In the text-to-SQL task, self-consistency is adapted to sampling a set of different SQL and voting for consistent SQL via execution feedback [30, 53]. Similarly, the SD+SA+Voting [52] framework eliminates those with execution errors identified by the deterministic database management system (DBMS) and opts for the prediction that garners the majority vote. Furthermore, motivated by recent research on extending the capabilities of LLMs with tools, FUXI [66] are proposed to enhance LLMs SQL generation through effectively invoking crafted tools.
$\mathbf{C}_{4}$-Execution Refinement
When designing criteria for accurate SQL generation, the priority is always whether a generated SQL can be successfully executed and retrieve content to correctly answer the user question [13]. As a complex programming task, generating the correct SQL in one go is challenging. Intuitively, considering the execution feedback/results in SQL generation assists the alignment to the corresponding database environment, which allows the LLMs to gather the potential executed errors and results to refine the generated SQL or hold a majority vote [30]. The execution-aware methods in text-to-SQL incorporate the execution feedback in two main approaches: 1) Incorporating the feedback through second round prompting for re-generation, for every SQL query generated in the initial response, it will be executed in the corresponding database, thus obtaining the database’s feedback. This feedback might be an error, or it might yield results that will be appended to the second round prompt. Through in-context learning of this feedback, LLMs are able to refine or re-generate the original SQL, thereby enhancing accuracy. 2) Utilize execution-based selection strategies for the generated SQL, sample multiple generated SQL queries from LLM, and execute each in the database. Based on the execution results of each SQL query, use selection strategies (e.g., self-consistency, majority vote [60]) to define one SQL query from the SQL set that satisfies the criteria as the final predicted SQL.
MRC-EXEC [67] introduced a natural language to code (NL2Code) translation framework with execution, which executes each sampled SQL query and selects the example with the minimal execution result–based Bayes risk [111]. LEVER [68] proposed an approach to verify NL2Code with execution, utilizing a generation and execution module to collect sampled SQL set and their execution results, respectively, then using a learned verifier to output the probability of the correctness. Similarly, the SELF-DEBUGGING [48] framework is presented to teach LLMs to debug their predicted SQL via few-shot demonstrations. The model is able to refine its mistakes by investigating the execution results and explaining the generated SQL in natural language without human interventions
As previously introduced, to incorporate the well-designed framework with execution feedback, two-stage implications are widely-used: 1. sampling a set of SQL queries. 2. majority vote (self-consistency). Specifically, the C3 [30] framework removes the errors and identifies the most consistent SQL; The retrieval-augmented framework [64] introduced a dynamic revision chain, combining fine-grained execution messages with database content to prompt the LLMs to convert the generated SQL query into natural language explanation; the LLMs are requested to identify the semantic gaps and revise their own generated SQL. Although schema-filtering approaches enhance SQL generation, the generated SQL could be unexecutable. DESEM [62] incorporates a fallback revision to address the issue; it revises and regenerates the SQL base on different kinds of errors and sets termination criteria to avoid the loop. DIN-SQL [8] designed a generic and gentle prompt in their self-correction module; the generic prompt requests the LLMs to identify and correct the error, and the gentle prompt asks the model to check the potential issue. The multi-agent framework MAC-SQL [57] comprises a refiner agent, which is able to detect and automatically rectify SQL errors, taking SQLite error and exception class to regenerate fixed SQL. Since different questions may require different numbers of revisions, SQL-CRAFT [55] framework introduced interactive correction with an automated control determination process to avoid over-correction or insufficient correction. FUXI [66] considers the error feedback in tool-based reasoning for SQL generation. The Knowledge-to-SQL [31] introduces a preference learning framework incorporating the database execution feedback with a direct preference optimization [112] for refining the proposed DELLM. PET-SQL [60] proposed cross consistency, which comprises two variants: 1) naive voting: instruct multiple LLMs to generate the SQL query, then utilizing the majority vote for the final SQL base on different execution results; 2) fine-grained voting: refine the naive voting based on the difficulty level to mitigate the voting bias.
@@ IV-B Fine-tuning
TABLE IV: Well-designed methods used in fine-tuning (FT) for LLM-based text-to-SQL. The methods in each category are ordered by release time. *The methods are utilized in multiple open-source LLMs; we select a representative model to present.
@@
CodeS [10] StarCoder [13, 33] Feb-2024 SIGMOD’24
StructLM [70] CodeLLaMA [13] Feb-2024 arXiv’24
Decomposition DTS-SQL [71] Mistral* [13, 40] Feb-2024 arXiv’24
Since supervised fine-tuning (SFT) is the mainstream approach in the LLMs training [29, 91], for open-source LLMs (e.g., LLaMA-2 [94], Gemma [113]), the most straightforward method to enable the model to adapt a specific domain quickly is to use collected domain label to perform SFT on the model. The SFT phase is typically the preliminary phase of the well-designed training framework [112, 114], as well as the fine-tuning of text-to-SQL. The auto-regressive generation process of SQL query $Y$ can be formulated as follows:
$P_{\pi}(Y\mid\mathcal{P})=\prod_{k=1}^{n}P_{\pi}(y_{k}\mid\mathcal{P},Y_{1:k-1}),$ (6)
where $Y=\{y_{1},y_{2},\ldots,y_{n}\}$ is an SQL query of length $n$, $y_{k}$ is the corresponding $k^{th}$ token of the SQL query, $Y_{1:k-1}$ is the prefix sequence of $Y$ ahead the token $y_{k}$. $P_{\pi}(y_{k}\mid\cdot)$ is a conditional probability of a LLM $\pi$ for generating the $k^{th}$ token of $Y$ base on the input prompt $\mathcal{P}$ and the prefix sequence.
Given a basic open-source model $\pi^{0}$, the goal of SFT is obtain a model $\pi^{SFT}$ through minimizing the cross-entropy loss:
$\mathcal{L}_{SFT}=-\sum_{k=1}^{n}\operatorname{log}P_{\pi^{0}}(\hat{y}_{k}=y_{k}\mid\mathcal{P},Y_{1:k-1}),$ (7)
where $\hat{y}_{k}$ is the $k^{th}$ token of the generated SQL query $\hat{Y}$, and $Y$ is the corresponding ground-truth label.
The SFT approach, which is also a vanilla fine-tuning approach for text-to-SQL has been widely adopted in text-to-SQL research for various open-source LLMs [10, 46, 9]. Compared to in-context learning (ICL) approaches, fine-tuning paradigms are more inclined to be at a starting point in LLM-based text-to-SQL. Currently, several studies exploring a better fine-tuning method have been released. We categorize the well-designed fine-tuning methods in different groups based on their mechanisms, as shown in Tab. IV.
Enhanced Architecture
The widely-used generative pre-trained transformer (GPT) framework utilizes decoder-only transformer architecture and conventional auto-regressive decoding for text generation. Recent studies on the efficiency of LLMs have revealed a common challenge: when generating long sequences with the auto-regressive paradigm, the need to incorporate the attention mechanism results in high latency for LLMs [115, 116]. In LLM-based text-to-SQL, the speed of generating SQL queries is significantly slower compared to traditional language modeling [21, 28], which has become a challenge in constructing high-efficiency local NLIDB.
As one of the solutions, CLLMs [69] are designed to address the above challenge with an enhanced model architecture and achieve a speedup for SQL generation.
Data Augmentation
During the fine-tuning process, the most straightforward factor affecting the model’s performance is the quality of the training labels [117]. The fine-tuning under the low quality or lack of the training labels is “making bricks without straw”, using high-quality or augmented data for fine-tuning always surpasses the meticulous design of fine-tuning methods on low-quality or raw data [74, 29]. Data-augmented fine-tuning in text-to-SQL made substantial progress, focusing on enhancing the data quality during the SFT process.
DAIL-SQL [9] are designed as an in-context learning framework, utilizing a sampling strategy for better few-shot instances. Incorporating the sampled instances in the SFT process improves the performance of open-source LLMs. Symbol-LLM [50] propose injection and infusion stage for data augmented instruction tuning. CodeS [10] augmented the training data with bi-directional generation with the help of ChatGPT. StructLM [70] are trained on multiple struct knowledge tasks for improving overall capability.
Pre-training

### v4

@@ IV-A1 $\mathbf{C}_{0}$-Vanilla Prompting
The impact of cross-domain examples is also being studied [66]. When incorporating in- and out-of-domain examples in different numbers, the in-domain demonstrations outperform zero-shot and out-of-domain examples, with performance improving as the number of examples increases. To explore the detailed construction of prompts, [73] compares concise and verbose prompt design strategies. The former style separates the schema, column names, primary keys, and foreign keys, while the latter organizes them as natural language descriptions.
IV-A2 $\mathbf{C}_{1}$-Decomposition
Decomposition-based methods reduce text-to-SQL complexity by dividing tasks or questions into manageable components, employing two main strategies:
1) Multi-module collaboration: This strategy breaks the task into specialized stages (e.g., schema linking, SQL generation, and SQL refinement) handled by distinct modules or agents. C3 [28] employs a three-module pipeline involving schema linking via clear prompting, calibration bias prompting for SQL guidance, and consistency-based voting for final selection. Similarly, MAC-SQL [48] and CHESS [56] use multi-agent frameworks where dedicated agents handle schema selection, decomposition, and refinement. Workflow-centric methods, like Distillery [58] and DEA-SQL [49], decompose the task into retrieval, generation, and correction stages. Iterative refinement and structural decomposition are also prevalent. DIN-SQL [7] integrates a four-stage pipeline, including schema linking, classification, SQL generation, and self-correction. Coder-Reviewer [46] separates generation and validation into distinct models. PET-SQL [52] refines outputs through multi-stage prompting, while MCS-SQL [55] uses execution-guided selection as the final module after multiple SQL generation.
2) Question decomposition: Complexity is reduced by splitting user questions into intermediate sub-questions via question decomposition. QDecomp [47] uses Least-to-Most prompting, breaking complex questions into step-by-step sub-questions to facilitate SQL generation. SGU-SQL [30] employs graph-based structure linking and meta-operators for decomposing grammar-aware sub-questions. Metadata-driven approaches, such as MetaSQL [50] and TA-SQL [54], further decompose questions using metadata conditioning and task-aligned schema linking. Meanwhile, frameworks like DIN-SQL [7] and MAC-SQL [48] use question decomposition as auxiliary modules but primarily focus on module-level collaboration.
IV-A3 $\mathbf{C}_{2}$-Prompt Optimization
Prompt optimization methods for text-to-SQL focus on enhancing prompt engineering quality. These methods are categorized into three key strategies:
1) Advanced few-shot sampling: This strategy optimizes example selection through semantic similarity, diversity, or domain alignment. DESEM [62] and retrieval-augmented frameworks [65, 70, 58] de-semanticize questions to retrieve examples with matching intent skeletons or vector embeddings. DAIL-SQL [8] balances quality and quantity by masking domain-specific tokens and ranking examples based on the Euclidean distance between questions and SQL queries, while SAFE-SQL [72] combines skeleton masking with self-augmentation. ODIS [66] hybridizes out-of-domain and synthetic in-domain examples for robustness, and FUSED [68] enhances diversity through iterative demonstration synthesis. Dynamic selection is explored in ACT-SQL [67] and MCS-SQL [55], which adaptively retrieve examples based on defined similarity scores or masked similarity scores.
2) Schema augmentation: These methods refine schema representations to reduce redundancy and improve relevance. C3 [28] distills question-relevant schemas by removing unrelated tables or columns, while RSL-SQL [59] simplifies schemas using NL descriptions. ReFoRCE [61] employs table compression and column exploration to manage long-context schema complexity. Gen-SQL [60] bridges NL-SQL gaps by generating pseudo-schemas that align with user questions. QDecomp [47] and SD+SA+Voting [64] further augment prompts with schema-aware semantic and structural hints.
3) External knowledge integration: This group enhances prompts with domain-specific knowledge or feedback mechanisms. Knowledge-to-SQL [69] trains a DELLM to inject external knowledge about question-database relationships, while ICRL [71] integrates reinforcement learning with schema-graph-derived knowledge bases. MAC-SQL [48] and DIN-SQL [7] implicitly utilize decomposition patterns as procedural knowledge to guide generation. C3 [28] and SAFE-SQL [72] incorporate calibration bias and self-augmented feedback loops, respectively, to iteratively refine outputs.
IV-A4 $\mathbf{C}_{3}$-Reasoning Enhancement
Reasoning enhancement methods [^4] Unlike decomposition paradigms, these methods retain single-turn generation while enhancing intrinsic reasoning through structured prompting. improve LLMs’ ability to handle complex SQL through structured reasoning and consistency mechanisms. These approaches focus on three crucial strategies:
1) Chain-of-thought (CoT)-based reasoning: CoT-based methods guide LLMs to generate intermediate reasoning steps before predicting SQLs. While standard CoT prompts (e.g., “Let’s think step by step”) show limited effectiveness in text-to-SQL [8, 1], adaptations help bridge this gap. ACT-SQL [67] automates CoT example generation by linking question slices to database columns via similarity matching. QDecomp [47] restructures SQL clauses into logical steps ordered by execution flow, reducing error propagation compared to generic CoT. CHASE-SQL [76] extends this with multi-path reasoning for candidate generation using “divide and conquer” CoT strategy.
2) Consistency-driven reasoning: These methods improve reliability through candidate diversity and agreement metrics before execution. Self-consistency [125] is employed in frameworks like C3 [28], DAIL-SQL [8], and SuperSQL [57], where majority voting selects the most frequent SQL candidate based on syntax or structural similarity. SD+SA+Voting [64] refines this by filtering candidates using schema-aware rules (e.g., valid joins, clause ordering) before voting. Unlike execution-dependent methods, these strategies prioritize logical consistency and syntactic validity over runtime feedback, making them lightweight but less robust to semantic errors.
3) Program-based and tool-augmented reasoning: Hybrid techniques combine SQL generation with external computation or tools. SQL-CRAFT [74] employs Program-of-Thoughts [126] (PoT) prompting, requiring LLMs to generate Python code alongside SQL for arithmetic-heavy queries, improving performance on datasets like BIRD [1]. FUXI [75] integrates dedicated tools for schema linking and syntax checking, dynamically invoking them during generation. These approaches differ from pure prompting strategies by explicitly bridging LLM reasoning with external symbolic systems.
IV-A5 $\mathbf{C}_{4}$-Execution Refinement
Execution refinement methods improve SQL accuracy by leveraging database feedback through regeneration or preferred SQL selection. These approaches are categorized into two paradigms:
1) Feedback-driven regeneration: These methods iteratively refine SQL queries using execution errors or intermediate results. Self-Debugging [79] enables LLMs to autonomously correct errors via natural language explanations of failed executions, while DESEM [62] uses error-type-specific prompts with termination criteria to prevent infinite loops. Multi-agent frameworks like MAC-SQL [48] and CHASE-SQL [76] utilize agents (e.g., Refiner, Query Fixer) to diagnose SQLite errors and regenerate SQL fixes. SQL-CRAFT [74] introduces adaptive control to balance correction depth, preventing over-refinement. Schema-focused methods like RSL-SQL [59] and DIN-SQL [7] use generic or gentle prompts to help LLMs identify schema misalignments during regeneration. Knowledge-to-SQL [69] optimizes its DELLM model through preference learning from execution feedback, while FUXI [75] integrates error feedback into tool-based reasoning. E-SQL [80] enhances questions with execution-aware candidate predictions.
2) Execution-guided selection: This strategy leverages database execution results to select or rank candidates. ReFoRCE [61] and PET-SQL [52] prune invalid SQL queries that fail execution, then apply bias-aware voting (e.g., difficulty-weighted or efficiency-prioritized as in MCS-SQL [55]). Probabilistic verification distinguishes methods: MRC-EXEC [77] minimizes Bayes risk by comparing execution outputs to expected results, and LEVER [78] trains a verifier to predict correctness likelihoods from execution traces. Hybrid approaches like the retrieval-augmented framework [65] use execution feedback for iterative SQL revision through semantic gap analysis, bridging generation and selection.
@@ IV-B Fine-tuning
TABLE IV: Well-designed methods used in the FT paradigm for LLM-based text-to-SQL. The methods in each category are ordered by release time. *The methods are utilized in multiple open-source LLMs; we select a representative model to present.
@@
CodeS [24] StarCoder [11, 1, 42] Feb-2024 SIGMOD’24
StructLM [83] CodeLLaMA [11] Feb-2024 COLM’24
Dubo-SQL [70] ChatGPT [1] Apr-2024 arXiv’24
Distillery [58] LLaMA-3.1* [1] Aug-2024 NeurIPS’24 Workshop
XiYan-SQL [84] Qwen2.5-Coder [11, 1] Nov-2024 arXiv’24
Multi-task Tuning SQL-LLaMA [48] CodeLLaMA [11, 1] Dec-2023 COLING’25
DTS-SQL [86] Mistral* [11, 37] Feb-2024 EMNLP’24 Findings
KaSLA [29] DeepSeek-Coder* [11, 1] Oct-2024 arXiv’24
MSc-SQL [87] Gemma-2 [11, 1] Oct-2024 NAACL’25
XiYan-SQL [84] Qwen2.5-Coder [11, 1] Nov-2024 arXiv’24
ROUTE [88] Qwen2.5* [11, 1, 37, 36, 38] Feb-2025 ICLR’25
Supervised Fine-tuning (SFT) is the widely-used approach within the FT paradigm for training LLMs [27, 111]. For open-source LLMs, like LLaMA [115] and Qwen [116] series, the most straightforward way to enable the model to adapt to a specific domain is to apply SFT over collected domain-specific data. The SFT phase typically serves as an initial step in a well-designed training pipeline [127, 128], including LLM-based text-to-SQL training. Consider the prompt defined in Eq. 4 and 5 as $P$, and let $Y=[y_{1},y_{2},\dots,y_{T}]$ represent the target ground-truth SQL query. The probability Pr for an LLM $\pi$ to generate $Y$ given $P$ is formally modeled as:
$\text{Pr}_{\pi}\left(Y\mid P\right)=\prod_{t=1}^{T}\text{Pr}_{\pi}\left(y_{t}\mid P,y_{<t}\right).$ (6)
Contrary to the LLM in the ICL paradigm, the parameters of $\pi$ are tuned. The SFT process minimizes the cross-entropy loss over the training dataset $\mathcal{D}$ for $\pi$:
$\mathcal{L}_{\text{SFT}}=-\sum_{\left(P,\,Y\right)\in\mathcal{D}}\sum_{t=1}^{T}\log\text{Pr}_{\pi}\left(y_{t}\mid P,y_{<t}\right).$ (7)
As a vanilla fine-tuning approach for LLM-based text-to-SQL, SFT is widely adopted in text-to-SQL research across various open-source LLMs [24, 119, 8]. FT paradigms are often more foundational in LLM-based text-to-SQL than ICL, which is the mainstream approach in the community. This is largely due to the comparably limited capabilities of open-source LLMs versus proprietary models in understanding complex database content. For the existing well-designed methods in the FT paradigm, we group these methods into different categories based on their training mechanisms, as shown in Tab. IV.
IV-B1 Enhanced Architecture
The widely-used Generative Pre-trained Transformer (GPT) framework [103] employs a decoder-only transformer architecture and conventional auto-regressive decoding for text generation. With advancements in the GPT framework, enhancements in LLM-based text-to-SQL involve modifications of the standard transformer backbone. These modifications incorporate tailored architecture to efficiently handle the structural and syntactic complexities of SQL. In traditional text-to-SQL research, this tailored architecture has effectively enhanced PLM-based performance. Techniques include schema-aware encoding [20], optimized attention [129], and targeted decoding [130], often relying on graph structures. Most recently, the speed of generating SQL queries by LLMs is notably slower than traditional PLM-based methods [19, 26], posing a challenge to building high-efficiency local NLIDBs.
As a solution, CLLMs [81] are designed to address this challenge with specialized decoding, offering speedup efficiency for SQL generation and reducing high latency.
IV-B2 Pre-training